OpenStudy (anonymous):

what is the x-intercept of the perpendicular line to y-3=3(x+1)

OpenStudy (anonymous):

\[y-3=3(x+1)\]\[y-3=3x+3\]\[y=3x+6\]the perpendicular slope is "negative multiplicative inverse" so, \[y= \frac{1}{3}x+6\]to find x-intercept make y=0. \[0=\frac{1}{3}x+6\]\[-\frac{1}{3}x=6\]\[x=-18\]

OpenStudy (anonymous):

your awesome thanks :)

OpenStudy (anonymous):

yw!

OpenStudy (anonymous):